# Zeroth-Order Policy Gradient for Reinforcement Learning from Human Feedback without Reward Inference

### v1

@@ Abstract
Reward inference (learning a reward model from human preferences) is a critical intermediate step in Reinforcement Learning from Human Feedback (RLHF) for fine-tuning Large Language Models (LLMs) such as ChatGPT. In practice, reward inference faces several fundamental challenges, including double problem misspecification, reward model evaluation without ground truth, distribution shift, and overfitting in joint reward model and policy training. An alternative approach that avoids these pitfalls is direct policy optimization without reward inference, such as Direct Preference Optimization (DPO), which provides a much simpler pipeline and has shown empirical success in LLMs. However, DPO utilizes the closed-form expression between the optimal policy and the reward function, which only works under the bandit setting or deterministic MDPs. This paper develops two RLHF algorithms without reward inference, which work for general RL problems beyond bandits and deterministic MDPs, and general preference models beyond the Bradely-Terry model. The key idea is to estimate the local value function difference from human preferences and then approximate the policy gradient with a zeroth-order gradient approximator. For both algorithms, we establish rates of convergence in terms of the number of policy gradient iterations, as well as the number of trajectory samples and human preference queries per iteration. Our results show there exist provably efficient methods to solve general RLHF problems without reward inference.
@@ 1 Introduction
In the past decade, we have witnessed unprecedented success in applying Reinforcement Learning (RL) to many applications, such as gaming AI (Knox and Stone, 2008; Warnell et al., 2018), recommendation and search (Zeng et al., 2016; Kohli et al., 2013), autonomous driving (Kiran et al., 2022) , and large language models (LLM) (Christiano et al., 2017; Wu et al., 2021; Nakano et al., 2021; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). RL studies the interaction between decision-making agents and a dynamic environment that keeps evolving. At each time step, the agent takes a certain decision (action) given the current state, a reward signal to measure the quality of that decision is provided by the environment. The agent’s goal is to learn a policy to maximize the cumulative reward, and the quality of the learned policy will depend on the per-step reward function. In classic RL, this reward function is usually handcrafted by domain experts to ensure it aligns with our expectations of the agent’s behavior. However, identifying a “good” reward function, also called Inverse RL (IRL), is one of the most fundamental problems in RL and is non-trivial. In recent years, researchers have developed a new approach called Reinforcement Learning from Human Feedback (RLHF) that uses human preference feedback as a signal to recover a reward function and then train or fine-tune LLMs (such as ChatGPT) with RL, which has delivered significant success. RLHF follows the diagram shown in Fig. 1, which includes three major steps (Ouyang et al., 2022): (i) policy network pre-training, (ii) query human evaluators for preferences over trajectories to train a reward model that aligns with human feedback, and (iii) use policy-gradient RL algorithms such as PPO to fine-tune the policy network using the reward model.
@@
Drawbacks of Reward Inference. To train a good reward model, i.e., to infer the underlying per-step reward function from human feedback (Christiano et al., 2017; Wang et al., 2023), the most common approach is to assume the feedback is generated based on a preference model such as the Bradley-Terry model (Bradley and Terry, 1952), and then maximize the log-likelihood of the collected trajectory comparison dataset accordingly over all possible (parameterized) reward functions. This procedure is indeed analyzed in most theoretical RLHF papers for both offline (Zhu et al., 2023; Zhan et al., 2024a) and online settings (Saha et al., 2023; Zhan et al., 2024b; Wu and Sun, 2024; Wang et al., 2023; Du et al., 2024). However, several challenges occur in practice for reward model training such as double problem mis-specification, reward model evaluation without ground truth, distribution shift, and overfitting in joint reward model and policy training (Casper et al., 2023). These drawbacks are also reflected in the theoretical results, e.g., overfitting of MLE in (Zhu et al., 2024). Moreover, similar to inverse reinforcement learning (IRL), the reward function that could explain the human feedback often is not unique, especially when given a limited amount of training trajectories (Arora and Doshi, 2021; Ng and Russell, 2000). Some reward models may make it difficult for agents to learn a good RL policy.
@@
Figure 1: A diagram illustrating classic policy-based RLHF and DPO: classic RLHF involves three steps: (i) pre-train a policy network (agent), (ii) collect trajectories from the environment using a behavior policy, query the human comparison for each trajectory pair and train a reward neural network through maximizing the likelihood under the Bradley-Terry model, and (iii) train the policy network with reward signals from the reward network. DPO does not train a reward network but directly optimizes the policy network from human preferences.
@@
DPO. To avoid the drawbacks of the reward inference in RLHF, Rafailov et al. (2023) proposed an algorithm called Direct Preference Optimization (DPO) which fine-tunes the LLM model directly from human preferences. Based on the Bradley-Terry preference model and a closed-form expression of the optimal policy given a reference policy and the reward function, DPO constructs a loss function directly from human feedback for learning the optimal policy, so avoids reward inference. This provides a much simpler pipeline and has great empirical performance (Rafailov et al., 2023, 2024a, 2024b). However, the closed-form expression of the optimal policy that DPO builds on is only for non-parametric policies and its theoretical justification only works for the bandit setting (Rafailov et al., 2023) or RL problems with deterministic transitions (Rafailov et al., 2024b). It remains an open question how to solve general RLHF problems without reward inference.
@@
RLHF without Reward Inference. Recently, value-based RLHF algorithms without global reward inference have been theoretically developed and analyzed (Xu et al., 2020; Zhang et al., 2024a) based on a dueling bandit approach (Bengs et al., 2021). The results, however, only hold for MDPs in tabular settings with finite state and action spaces. Chen et al. (2022) studied the function approximation regime, but their algorithm requires both the true preference model and the transition kernel to belong to a known function class, which is impractical. The result also depends on the function class complexity which is usually large for most function approximators in practice. So far, no provable policy-based algorithm in this category has been developed.
@@ 1.1 Main Contributions
where $r(x,a)$ is the reward in state $x$ with action $a,$ $\pi_{\hbox{\scriptsize ref}}$ is a reference policy and $\pi^{*}$ is the optimal policy. Based on the direction connection, the policy optimization can be formulated as a direct matching between human preference and the optimal policy with a log-likelihood loss function. In a recent paper (Rafailov et al., 2024a), it has been further shown that DPO solves a KL-divergence-constrained policy optimization problem for the deterministic token-level MDP for LLM, where the next state is deterministic given the current state and action. For general RL problems with parameterized policies, equation 1 does not hold, and it is often hard if not impossible to obtain a “global” function like it that connects the optimal policy and the reward (hence human feedback).
This paper exploits the “local” relation between human feedback and policy optimization. In particular, given a policy $\pi_{\bm{\theta}}$ and a perturbed version of the policy $\pi_{{\bm{\theta}}+v}$, we use human feedback over the trajectories generated from both policies to inform the more preferred policy. Intuitively, if one trajectory is preferred over the other, the policy that generates this trajectory is likely to have a higher value. Then given a preference model such as the Bradley-Terry model, we can further estimate the value function differences of the two policies, $V(\pi_{{\bm{\theta}}+v})-V(\pi_{\bm{\theta}})$, using the population-level preference, where $V(\pi)$ is the value function associated with policy $\pi.$ Finally, the value difference can be used as an estimator of policy gradient, $\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}})$, following the zeroth-order optimization approach (Nesterov and Spokoiny, 2017; Ghadimi and Lan, 2013) to improve the policy.
Based on this idea, this paper proposes two RLHF algorithms without reward inference: Zeroth-Order Policy Gradient (ZPG) and Zeroth-Order Block-Coordinate Policy Gradient (ZBCPG), both from Human Feedback. ZBCPG differs from ZPG in its policy perturbation rule, which has lower computational complexity and allows parallel optimization since one can sample multiple perturbed policies to perform policy gradient and combine the estimated gradient. Under mild assumptions, both algorithms have the following rate of converge to a stationary point:
@@
We remark that (Tang et al., 2024a) proposes a similar approach towards utilizing human feedback and a zeroth-order gradient descent algorithm from ranking data. However, they assume an error-free ranking oracle over policies based on their value functions, which makes their problem a deterministic optimization problem and does not apply to trajectory preference data like in RLHF and DPO. This paper studies RLHF with trajectory preferences and quantifies the impacts of stochastic trajectories and human preferences on the rate of convergence of RLHF without reward inference.
2 Related Work
@@
In this section, we review recent developments of RLHF and zeroth-order optimization for both empirical and theoretical results.
@@
Empirical Studies of RLDHF in LLMs. DPO (Rafailov et al., 2024a) and SLiC-HF (Zhao et al., 2023) have empirically shown it is possible to directly learn an RL policy from human preference. In DPO, the authors solve a KL-divergence-constrained reward maximization problem for a prompt response generation problem similar to the contextual bandit. The optimal policy for this problem has a closed-form expression, and the reward of each prompt-response pair can be computed knowing the optimal policy and the reference policy. The authors then minimize a log-likelihood loss by plugging the reward expression into the Bradley-Terry model to measure the alignment between the policy and the human preference without training a reward network. Rafailov et al. (2024b) extends this loss to token-level MDP on the condition that the MDP transition is deterministic, i.e., the next state is a concatenation between the current state and action. This limits the DPO loss minimization approach to LLM problems. In general MDPs with stochastic transition, the DPO loss cannot be computed following the derivation in (Rafailov et al., 2024b)Azar et al. (2024) extends DPO to a wider class of RL problems, even avoiding the notion of an underlying reward function. Instead of maximizing the reward in a KL-constrained problem like DPO, the authors proposed to optimize a general non-decreasing function of the ground-truth population-level preference probability. Other variants are also studied. Ethayarajh et al. (2024) considers aligning policy with humans and designs loss from a prospect theory perspective, and Tang et al. (2024b) considers optimizing a general loss of the preference loss instead of the log-likelihood. Dong et al. (2024) and Xiong et al. (2024) proposed to obtain human feedback in an online fashion to mitigate the distribution-shift and over-parameterization phenomenon. Attempts to understand the theoretical performance of RLDHF algorithms such as DPO are made in (Azar et al., 2024), but the authors only showed the existence of optima of the loss function, without any policy optimality and sample complexity guarantees.
@@
RLHF with Provable Sample Complexity. Two major approaches have been studied to learn the optimal policy from human preference data. The first is similar to the traditional RLHF paradigm such as PPO used in empirical studies which infers a reward function, or sometimes a utility function, and then trains an RL agent. This reward inference step is also called surrogate learning in the preference-based RL literature, i.e., see (Wirth et al., 2017, Sec. 2.4). The second approach directly optimizes the policy from human preferences. Empirical algorithms for both approaches have been developed without theoretical guarantees for a few years. For example, reward function or utility function estimation followed by an RL policy search algorithm has been proposed and studied in (Schoenauer et al., 2014; Wirth et al., 2016; Christiano et al., 2017), while direct policy learning from humans through trajectory sampling also has been proposed in using a heuristic evolutionary algorithm (Busa-Fekete et al., 2014; Akrour et al., 2011), or from a Bayesian Markov chain Monte Carlo perspective (Wilson et al., 2012). However, it was not until recently that algorithms with provable theoretical guarantees are proposed. From the reward inference approach, Novoseller et al. (2020) took a Bayesian perspective and maintains a posterior distribution over the reward function and the transition kernel, which is computationally costly. Wang et al. (2023) assumes the reward function can be linearly parameterized with known feature embedding and proposes a preference-to-reward interface (a reward model) using an online policy update and a baseline trajectory for comparison. The authors theoretically showed that for a general known preference model setting, RLHF is no harder than RL. A similar analysis framework is adopted in contemporary theoretical works, e.g., (Saha et al., 2023; Zhan et al., 2024a, b; Zhu et al., 2023; Kong and Yang, 2022; Wu and Sun, 2024), for both online, offline, and hybrid RL problems, where the algorithms first learn the linear parameter of the reward function and perform value-based RL on the learned reward function. The analysis first characterizes the error of the reward parameter using the concentration of the MLE estimator and then propagates this error to the value-based RL algorithm. Specifically, Saha et al. (2023) considers the tabular RL setting and assumes a known feature embedding for each trajectory with a known transition kernel. Then, the agent is directed by RL to explore the feature direction where the uncertainty in the reward function is large. Wu and Sun (2024) extends the scenario to linear MDPs with unknown transition and uses least square value iteration to solve the RL problem. Zhu et al. (2023) extends the work (Saha et al., 2023) to the offline RL scenario where a pessimistic estimator of the reward parameter is used to combat the insufficient data coverage in offline settings. Zhan et al. (2024b) replaces the linear reward parameterization with a general function class under the realizability condition. This also enables them to solve the unknown preference model as long as it is in the known function class. Zhan et al. (2024b) and Kong and Yang (2022) studies the hybrid RL problem with human preference, where they first use an exploratory policy from an optimal design perspective to improve the coverage of the offline dataset and to extract more information useful for reward inference. Then, the problem will be solved under the general RL framework. All papers above took a value-based approach in RL and Du et al. (2024) analyzed natural policy gradient with reward inference. Direct policy learning from human preference has been less analyzed compared to reward inference approaches. In tabular MDPs, Xu et al. (2020) and Zhang et al. (2024a) reduce the RL problem to a sequence of dueling bandit problems. However, the approach is only suitable for MDPs with finite state and action spaces. For general MDPs, Chen et al. (2022) first uses function approximation for the mapping from trajectory pair to human preference, and learns the RL transition kernel from a least square estimator with this preference approximator. The optimal policy can then be learned using a dueling bandit approach. However, their results assume the true preference model and transition kernel is inside the known function class with small complexity, which is strong in real applications. RLHF has also been studied in other aspects. Li et al. (2024) studies RL from human behavior dataset from a dynamic choice perspective.  Chakraborty et al. (2024) formulated the reward inference and the policy optimization as a bilevel optimization problem, and Kausik et al. (2024) studies RLHF with partially observed rewards and states. Zhu et al. (2024) studies the overfitting issue in reward model training.
@@
Zeroth-Order Optimization. The zeroth-order optimization problem has been studied in the convex and non-convex optimization literature for more than a decade (Ghadimi and Lan, 2013; Nesterov and Spokoiny, 2017), where the stochastic gradient descent algorithm with two-point gradient estimator is most widely used. The convergence rate in smooth functions is first studied in (Ghadimi and Lan, 2013) in both convex and non-convex settings. In the convex setting, the algorithm finds the optimal point while in the non-convex setting, the algorithm finds a stationary point. The rate of convergence for non-convex functions is improved in Nesterov and Spokoiny (2017). Variants of stochastic gradient descent such as variance reduction techniques (Liu et al., 2018a) and ADMM (Liu et al., 2018b; Gao et al., 2018) have also been studied in the zeroth-order literature. (Cai et al., 2021) extends the zeroth-order method to blocked coordinate descent for computational efficiency and Liu et al. (2019) extends the method to analyze a signed version of SGD which is more memory efficient in federated settings since each element of the gradient takes only one bit. Recently, the zeroth-order optimization technique has been proposed in optimizing LLMs (Malladi et al., 2023; Zhang et al., 2024b), but they only implement it in the procedure of policy optimization from the reward network to avoid the heavy memory burden in back-propagation. Our work is different in that we view human feedback as a natural source of zeroth-order information and apply the method directly from human preference.

### v2

@@ Abstract
Reward inference (learning a reward model from human preferences) is a critical intermediate step in the Reinforcement Learning from Human Feedback (RLHF) pipeline for fine-tuning Large Language Models (LLMs). In practice, RLHF faces fundamental challenges such as distribution shift, reward model overfitting, and problem misspecification. An alternative approach is direct policy optimization without reward inference, such as Direct Preference Optimization (DPO), which provides a much simpler pipeline and has shown empirical success in LLM applications. However, DPO utilizes the closed-form expression between the optimal policy and the reward function, which is only suitable under the bandit setting or deterministic MDPs. This paper develops two RLHF algorithms without reward inference for general RL problems beyond bandits and deterministic MDPs, and general preference models beyond the Bradley-Terry model. The key idea is to estimate the local value function difference from human preferences and then approximate the policy gradient with a zeroth-order gradient approximator. For both algorithms, we establish polynomial convergence rates in terms of the number of policy gradient iterations, the number of trajectory samples, and human preference queries per iteration. Numerical experiments in stochastic environments validate the performance of our proposed algorithms, outperforming popular RLHF baselines such as DPO and PPO. Our paper shows there exist provably efficient methods to solve general RLHF problems without reward inference.
@@ 1 Introduction
In the past decade, we have witnessed unprecedented success in applying Reinforcement Learning (RL) to many applications, such as video games (Knox and Stone, 2008; Warnell et al., 2018), recommendation and search (Zeng et al., 2016; Kohli et al., 2013), autonomous driving (Kiran et al., 2022). RL studies the interaction between decision-making agents and an evolving dynamic environment. At each time step, the agent takes a certain decision (action) given the current state, a reward signal to measure the quality of that decision is provided by the environment. The agent’s goal is to learn a policy to maximize the cumulative reward, and the quality of the learned policy will depend on the per-step reward function. In classic RL, this reward function is usually handcrafted by domain experts to ensure it aligns with human interests. However, the problem of identifying a “good” reward function, also referred to as Inverse Reinforcement Learning (IRL), is non-trivial and one of the most fundamental problems in the history of RL (Ng and Russell, 2000). In recent years, Reinforcement Learning from Human Feedback (RLHF) that uses human preference feedback as a signal to recover a reward function has emerged to fine-tune Large Language Models (LLMs), which has delivered significant success (Christiano et al., 2017; Wu et al., 2021; Nakano et al., 2021; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). RLHF follows the diagram shown in Fig. 1, which includes three major steps (Ouyang et al., 2022): (i) pre-train a policy neural network, (ii) collect sets of trajectory pairs and query human evaluators for preferences over trajectory pairs to train a reward model using maximum likelihood to align with human feedback, and (iii) use policy-optimization-based RL algorithms such as PPO (Schulman et al., 2017) to fine-tune the policy network with the reward signals generated by the reward model. The reward inference intermediate step, which trains the reward network, is crucial to obtaining a high-quality policy through RL in the final step.
@@
Drawbacks of Reward Inference. To train a good reward model, i.e., to infer the underlying per-step reward function from human feedback (Christiano et al., 2017; Wang et al., 2023), the most common approach is to assume the feedback is generated based on a preference model such as the Bradley-Terry model (Bradley and Terry, 1952), and then maximize the log-likelihood of the collected trajectory comparison dataset accordingly over all possible (parameterized) reward functions. This procedure is indeed analyzed in most theoretical RLHF papers for both offline (Zhu et al., 2023; Zhan et al., 2024a) and online settings (Saha et al., 2023; Zhan et al., 2024b; Wu and Sun, 2024; Wang et al., 2023; Du et al., 2024). However, several challenges occur in practice for reward model training such as double problem misspecification, reward model evaluation without ground truth, distribution shift, and overfitting in joint reward model and policy training (Casper et al., 2023). These drawbacks are also reflected in the theoretical results, e.g., overfitting of the maximum likelihood estimator (MLE) in Zhu et al. (2024). Moreover, similar to the dilemma in IRL, the reward function that could explain human feedback is often not unique, especially when given limited training trajectories (Arora and Doshi, 2021; Ng and Russell, 2000). Some reward models may make it difficult for agents to learn a good RL policy.
@@
Figure 1: A diagram illustrating classic policy-based RLHF and DPO: classic RLHF involves three steps: (i) policy pre-training: pre-train a policy network (agent), (ii) reward inference: collect trajectories from the environment using a behavior policy, query the human comparison for each trajectory pair and train a reward neural network through maximizing the likelihood under the Bradley-Terry model, and (iii) policy training with reward model: train the policy network with reward signals sampled from the reward network. DPO does not train a reward network but directly optimizes the policy network from human preferences.
@@
DPO. To avoid the drawbacks of the reward inference in RLHF, Rafailov et al. (2023) proposed an algorithm called Direct Preference Optimization (DPO) which fine-tunes the LLM directly from human preferences. Based on the Bradley-Terry preference model and a closed-form expression of the optimal policy given a reference policy and the reward function, DPO constructs a loss function directly from human feedback for learning the optimal policy to avoid reward inference. This provides a much simpler pipeline and has great empirical performance (Rafailov et al., 2023, 2024a, 2024b). However, the closed-form expression of the optimal policy that DPO builds on is only for non-parametric policies and its theoretical justification only works for the bandit setting (Rafailov et al., 2023) or RL problems with deterministic transitions (Rafailov et al., 2024b). It remains an open question how to solve general RLHF problems without reward inference.
@@
RLHF without Reward Inference. Recently, value-based RLHF algorithms without global reward inference have been theoretically developed and analyzed (Xu et al., 2020; Zhang et al., 2024a) based on a dueling bandit approach (Bengs et al., 2021). The results, however, only hold for MDPs in tabular settings with finite state and action spaces. Chen et al. (2022) studied the function approximation regime, but their algorithm requires both the true preference model and the transition kernel to belong to a known function class, which is also far from practice. The result also depends on the function class complexity, which is usually large for most function approximators in practice. So far, no provable policy-based algorithm in this category has been developed.
@@ 1.1 Main Contributions
where $r(x,a)$ is the reward in state $x$ with action $a,$ $\pi_{\hbox{\scriptsize ref}}$ is a reference policy and $\pi^{*}$ is the optimal policy. Based on the direction connection, the policy optimization can be formulated as a direct matching between human preference and the optimal policy with a log-likelihood loss function. In a recent paper (Rafailov et al., 2024a), it has been further shown that DPO solves a KL-divergence-constrained policy optimization problem for the deterministic token-level MDP for LLMs, where the next state is deterministic given the current state and action. For general RL problems with parameterized policies, equation 1 does not hold, and it is often hard if not impossible to obtain a “global” function like it that connects the optimal policy and the reward (hence human feedback).
This paper exploits the “local” relation between human feedback and policy optimization. In particular, given a policy $\pi_{\bm{\theta}}$ and a perturbed version of the policy $\pi_{{\bm{\theta}}+{\bm{v}}}$ where ${\bm{v}}$ is a small perturbation vector, we use human feedback over the trajectories generated from both policies to inform the direction of a more preferred policy. Intuitively, if one trajectory is preferred over the other, the policy that generates this trajectory is likely to have a higher value. Then given a preference model such as the Bradley-Terry model, we can further estimate the value function differences of the two policies, $V(\pi_{{\bm{\theta}}+{\bm{v}}})-V(\pi_{\bm{\theta}})$, where $V(\pi)$ is the value function associated with policy $\pi.$ Finally, the value difference can be used as an estimator of policy gradient, $\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}})$, following the zeroth-order optimization approach (Nesterov and Spokoiny, 2017; Ghadimi and Lan, 2013) to improve the policy.
Based on this idea, this paper proposes two RLHF algorithms without reward inference: Zeroth-Order Policy Gradient (ZPG) and Zeroth-Order Block-Coordinate Policy Gradient (ZBCPG), both from Human Feedback. ZBCPG differs from ZPG in its policy perturbation rule, which has lower computational complexity and allows parallel optimization since one can sample multiple perturbed policies to perform policy gradient and aggregate the estimated gradient. Under mild assumptions, both algorithms have the following rate of convergence to a stationary policy:
@@
We remark that Tang et al. (2024a) proposes a similar approach towards utilizing human feedback and a zeroth-order gradient descent algorithm from ranking data. However, they assume an error-free ranking oracle over policies based on their value functions, which makes their problem a deterministic optimization problem and does not apply to trajectory preference data like in RLHF and DPO. This paper studies RLHF with trajectory preferences and quantifies the impacts of stochastic trajectories and human preferences on the rate of convergence of RLHF without reward inference.
1.2 Related Work
@@
We review recent developments in RLHF and zeroth-order optimization for both empirical and theoretical results. A thorough survey on RLHF can be found in Kaufmann et al. (2023) and Casper et al. (2023)
@@
Empirical Studies of Direct RLHF in LLMs. DPO (Rafailov et al., 2024a) and SLiC-HF (Zhao et al., 2023) have empirically shown it is possible to directly learn an RL policy from human preference. In DPO, the authors solve a KL-divergence-constrained reward maximization problem for a prompt response generation problem similar to the contextual bandit. The optimal policy for this problem has a closed-form expression, and the reward of each prompt-response pair can be computed knowing the optimal policy and the reference policy. The authors then minimize a log-likelihood loss by plugging the reward expression into the Bradley-Terry model to measure the alignment between the policy and the human preference without training a reward network. Rafailov et al. (2024b) extends this loss to token-level MDP on the condition that the MDP transition is deterministic, i.e., the next state is a concatenation between the current state and action. This limits the DPO loss minimization approach to LLM problems. In general MDPs with stochastic transition, the DPO loss cannot be computed following the derivation in (Rafailov et al., 2024b)Azar et al. (2024) extends DPO to a wider class of RL problems, even avoiding the notion of an underlying reward function. Instead of maximizing the reward in a KL-constrained problem like DPO, the authors proposed to optimize a general non-decreasing function of the ground-truth population-level preference probability. Other variants are also studied. Ethayarajh et al. (2024) considers aligning policy with humans and designs loss from a prospect theory perspective, and Tang et al. (2024b) considers optimizing a general loss of the preference loss instead of the log-likelihood. Dong et al. (2024) and Xiong et al. (2024) proposed to obtain human feedback in an online fashion to mitigate the distribution-shift and over-parameterization phenomenon. Attempts to understand the theoretical performance of RLDHF algorithms such as DPO are made in (Azar et al., 2024), but the authors only showed the existence of optima of the loss function, without any policy optimality and sample complexity guarantees.
@@
RLHF with Provable Sample Complexity. Two major approaches have been studied to learn the optimal policy from human preference data. The first is similar to the traditional RLHF paradigm, such as PPO used in empirical studies, which infers a reward function, or sometimes a utility function, and then trains an RL agent. This reward inference step is also called surrogate learning in the preference-based RL literature, i.e., see (Wirth et al., 2017, Sec. 2.4). The second approach directly optimizes the policy from human preferences. Empirical algorithms for both approaches have been developed without theoretical guarantees for a few years. For example, reward function or utility function estimation followed by an RL policy search algorithm has been proposed and studied in (Schoenauer et al., 2014; Wirth et al., 2016; Christiano et al., 2017), while direct policy learning from humans through trajectory sampling also has been proposed in using a heuristic evolutionary algorithm (Busa-Fekete et al., 2014; Akrour et al., 2011), or from a Bayesian Markov chain Monte Carlo perspective (Wilson et al., 2012). However, it was not until recently that algorithms with provable theoretical guarantees were proposed. From the reward inference approach, Novoseller et al. (2020) took a Bayesian perspective and maintained a posterior distribution over the reward function and the transition kernel, which is computationally costly. Wang et al. (2023) assumes the reward function can be linearly parameterized with known feature embedding and proposes a preference-to-reward interface (a reward model) using an online policy update and a baseline trajectory for comparison. The authors theoretically showed that for a general known preference model setting, RLHF is no harder than RL. A similar analysis framework is adopted in contemporary theoretical works, e.g., (Saha et al., 2023; Zhan et al., 2024a, b; Zhu et al., 2023; Kong and Yang, 2022; Wu and Sun, 2024), for both online, offline, and hybrid RL problems, where the algorithms first learn the linear parameter of the reward function and perform value-based RL on the learned reward function. The analysis first characterizes the error of the reward parameter using the concentration of the MLE estimator and then propagates this error to the value-based RL algorithm. Specifically, Saha et al. (2023) considers the tabular RL setting and assumes a known feature embedding for each trajectory with a known transition kernel. Then, the agent is directed by RL to explore the feature direction where the uncertainty in the reward function is large. Wu and Sun (2024) extends the scenario to linear MDPs with unknown transition and uses least square value iteration to solve the RL problem. Zhu et al. (2023) extends the work (Saha et al., 2023) to the offline RL scenario where a pessimistic estimator of the reward parameter is used to combat the insufficient data coverage in offline settings. Zhan et al. (2024b) replaces the linear reward parameterization with a general function class under the realizability condition. This also enables them to solve the unknown preference model as long as it is in the known function class. Zhan et al. (2024b) and Kong and Yang (2022) study the hybrid RL problem with human preference, where they first use an exploratory policy from an optimal design perspective to improve the coverage of the offline dataset and to extract more information useful for reward inference. Then, the problem will be solved under the general RL framework. All papers above took a value-based approach in RL and Du et al. (2024) analyzed natural policy gradient with reward inference. Direct policy learning from human preference has been less analyzed compared to reward inference approaches. In tabular MDPs, Xu et al. (2020) and Zhang et al. (2024a) reduce the RL problem to a sequence of dueling bandit problems. However, the approach is only suitable for MDPs with finite state and action spaces. For general MDPs, Chen et al. (2022) first uses function approximation for the mapping from trajectory pair to human preference, and learns the RL transition kernel from a least square estimator with this preference approximator. The optimal policy can then be learned using a dueling bandit approach. However, their results assume the true preference model and transition kernel are inside the known function class with small complexity, which is strong in real applications. RLHF has also been studied in other aspects. Li et al. (2023) studies RL from the human behavior dataset from a dynamic choice perspective.  Chakraborty et al. (2024) formulated the reward inference and the policy optimization as a bilevel optimization problem, and Kausik et al. (2024) studies RLHF with partially observed rewards and states. Zhu et al. (2024) studies the overfitting issue in reward model training. Recently, Xie et al. (2024) combined DPO with optimistic exploration to design XPO in the function approximation regime with provable convergence. However, their algorithm and results are still dependent on the DPO loss, which only works for deterministic MDPs.
@@
Zeroth-Order Optimization and Evolutionary Strategies. The zeroth-order optimization problem has been studied in the convex and non-convex optimization literature for more than a decade (Ghadimi and Lan, 2013; Nesterov and Spokoiny, 2017), where the stochastic gradient descent algorithm with two-point gradient estimator is most widely used. The convergence rate in smooth functions is first studied in (Ghadimi and Lan, 2013) in both convex and non-convex settings. In the convex setting, the algorithm finds the optimal point while in the non-convex setting, the algorithm finds a stationary point. The rate of convergence for non-convex functions is improved in Nesterov and Spokoiny (2017). Variants of stochastic gradient descent such as variance reduction techniques (Liu et al., 2018a) and ADMM (Liu et al., 2018b; Gao et al., 2018) have also been studied in the zeroth-order literature. (Cai et al., 2021) extends the zeroth-order method to blocked coordinate descent for computational efficiency and Liu et al. (2019) extends the method to analyze a signed version of SGD which is more memory efficient in federated settings since each element of the gradient takes only one bit. In recent years, the study of evolutionary strategies (Rechenberg, 1973) brought zeroth-order optimization methods to optimize RL agents, which has achieved empirical success as a scalable alternative to classic RL algorithms such as Q-learning or policy gradient (Salimans et al., 2017; Conti et al., 2018). This strategy has also been applied in preference-based RL as well (Busa-Fekete et al., 2014; Akrour et al., 2011). However, the theoretical guarantees and provable algorithms remain under-explored. The zeroth-order optimization technique has been proposed in optimizing LLMs (Malladi et al., 2023; Zhang et al., 2024b), but they only implement it in the procedure of policy optimization from the reward network to avoid the heavy memory burden in back-propagation. In their studied problem, the loss function can be explicitly estimated or calculated, and thus can be queried to construct the zeroth-order gradient, which is more similar to classic evolutionary strategies. Our work is different in that the loss function, i.e., the value function, cannot be directly queried without reward feedback, so the zeroth-order algorithms cannot be directly used. However, we view human feedback as a natural source of zeroth-order information and apply the method directly from human preference.
Preference Models. The study of modeling the rationale of human decision-making has been pertinent for almost a century. The most popular and widely studied model in the literature is the random utility model (RUM) in social choice theory (Azari et al., 2012), which was first developed as early as 1920s (Thurstone, 1927). In RUM, for each decision-maker, each choice is associated with a utility that consists of two portions: a shared public utility same for all people, and a private utility (noise) unique to each person. People are modeled as utility maximizers and therefore will choose the choice that has the largest total utility. The private utility is assumed to follow a certain distribution among the population of decision-makers. It is not hard to see that the public utility will determine which choice is more preferred among the population, and the distribution of the private utility will determine how much it is more preferred, i.e., the population preference probability. Different distribution assumptions give rise to different preference models with different link functions. If the private utility follows a normal distribution, the preference model is called the Probit model where the link function is the cumulative distribution function of a normal distribution. The Probit model is the first proposed and studied preference model in the literature (Thurstone, 1927). Suppose the private utility follows the Gumbel distribution. We recover the Bradley-Terry model, also known as the logit model, with the logistic link function. Other preference models include the linear model with linear link function, the Weibull model, the Cauchy model, and the complementary log-log model, which have been studied in the literature for specific applications, e.g., see (Train, 2009; Greene, 2000, 2010) for a comprehensive review. In general, when the utility is close among choices, the Bradley-Terry model and the Probit model are often the most accurate and easy to implement.